# Know You First and Be You Better: Modeling Human-Like User Simulators via Implicit Profiles

### v2

@@ Abstract
User simulators are crucial for replicating human interactions with dialogue systems, supporting both collaborative training and automatic evaluation, especially for large language models (LLMs). However, existing simulators often rely solely on text utterances, missing implicit user traits such as personality, speaking style, and goals. In contrast, persona-based methods lack generalizability, as they depend on predefined profiles of famous individuals or archetypes. To address these challenges, we propose User Simulator with implicit Profiles (USP), a framework that infers implicit user profiles from human-machine conversations and uses them to generate more personalized and realistic dialogues. We first develop an LLM-driven extractor with a comprehensive profile schema. Then, we refine the simulation through conditional supervised fine-tuning and reinforcement learning with cycle consistency, optimizing it at both the utterance and conversation levels. Finally, we adopt a diverse profile sampler to capture the distribution of real-world user profiles. Experimental results demonstrate that USP outperforms strong baselines in terms of authenticity and diversity while achieving comparable performance in consistency. Furthermore, dynamic multi-turn evaluations based on USP strongly align with mainstream benchmarks, demonstrating its effectiveness in real-world applications  [^1] All resources will be released at https://github.com/wangkevin02/USP.
@@
Kuang Wang1,2, Xianfei Li1, Shenghao Yang1, Li Zhou1, Feng Jiang1,2,3 thanks: Feng Jiang is the corresponding author., Haizhou Li1,2 1The Chinese University of Hong Kong, Shenzhen 2Shenzhen Research Institute of Big Data 3Shenzhen University of Advanced Technology kuangwang@link.cuhk.edu.cn, jeffreyjiang@cuhk.edu.cn
@@ 1 Introduction
Figure 1: Examples of performance across various user simulators in multi-turn human-LLM interactions. OF and SC refer to objective facts and subjective characteristics, respectively.
@@
The user simulator is designed as a proxy for real users in interactions with large language models (LLMs). It can simulate realistic user by generating the target user’s behavior or utterances based on the specified characteristics, enabling dynamic multi-turn interactions with LLMs Wan et al. (2022) and scene reproduction Wang et al. (2024b). As a result, it becomes an effective alternative Liu et al. (2023); Ferreira et al. (2024) in scenarios where real-world human-computer interaction data is difficult to obtain, especially in domains with privacy and ethical concerns, such as medical consultations Valizadeh and Parde (2022). It also helps Simulation-to-Reality (Sim2Real) applications, such as tutorial strategies, election simulations, and public opinion research Liu et al. (2024); Zhang et al. (2024b); Chuang et al. (2024).
Recent LLMs advance promote user simulators by enhancing naturalness and utility Deng et al. (2024); Zhang et al. (2024a), as shown in Figure 1. Since directly using LLM as the user simulator suffers from role confusion Xu et al. (2023a), some works Xu et al. (2023a); Kong et al. (2024); Sun et al. (2024) attempt to enhance its authenticity by training a user simulator on the conversation dataset. However, they are only trained on text utterance, making it difficult to simulate diverse user behaviors without seed context, exhibiting limited self-awareness Tseng et al. (2024) and failing to maintain a consistent personality. Although some LLM-based role-playing methods Moon et al. (2024) utilize predefined profiles to alleviate this problem, they require additional extensive annotations and can only be applied to celebrities, lacking the diversity of simulation.
@@
To address the issues above, we believe that a user simulator knows users’ intrinsic characters hidden in their conversations first and then can be a better simulation. Therefore, we treat user simulation as a dialogue reconstruction task and propose a novel framework named the User Simulator with implicit Profile (USP). It is decomposed into implicit profile extraction to capture the user’s underlying characteristics from the target user dialogue and conditional generation based on the profile.
@@
In this framework, we first propose an LLM-driven profile extractor to extract implicit profiles from user conversations with a well-designed profile schema. Inspired by interpersonal interaction theoryKruglanski and Higgins (2013), our profile schema contains two dimensions (objective facts (OF) and subjective characteristics (SC)) with a dozen attributes to describe the user comprehensively. Different from existing works Cheng et al. (2024b); Tu et al. (2024), we then polish the profile attributes into natural, descriptive profiles to ensure generalization.
@@
Our experiments reveal that USP enhances semantic and stylistic similarity of reconstructed multi-turn dialogues to target dialogues by about 34% and 43% over the leading baseline, with reconstruction errors halved, showcasing superior authenticity and diversity. It matches the consistency of ProfileGPT(4o), a GPT-4o-based simulation baseline, boosting dialogue profile consistency by 14% in multi-turn scenarios while performing comparably in single-turn cases. Furthermore, our USP-based multi-turn dynamic evaluation of LLMs for downstream tasks aligns well with established benchmarks,allowing for a more granular assessment of LLM performance across different user groups. Our key contributions are outlined below:
@@
  • We develop a new framework that infers implicit user profiles as insight, further enhanced with conditional fine-tuning and reinforcement learning with cycle consistency for better simulation at both the utterance and conversation levels.
  • Experimental results show USP surpasses all baselines in authenticity and diversity, maintains comparable consistency, and proves effective for multi-turn dynamic evaluation of LLMs.
@@ 2.1 General User Simulator
Early user simulators including agenda-based methods Schatzmann et al. (2007); Schatzmann and Young (2009) and model-based methods  Asri et al. (2016); Kreyssig et al. (2018). These simulators were initially designed with a narrow scope due to limited natural language generation capabilities, such as generating synthetic binary preference responses Christakopoulou et al. (2016) in conversational recommendation systems.
@@ 2.2 Persona-based User Simulator
General user simulators often struggle to capture the full spectrum of diverse user needs, leading to a growing interest in persona-based personalization to improve both controllability and diversity in simulations Takanobu et al. (2020). Some researchers attempt to leverage goal generators Takanobu et al. (2020) to create diverse user goals or retrieval-based personas derived from historical data Shi et al. (2019) to guide user simulators in task-oriented dialogue (ToD) systems.
With the rise of LLMs and their impressive zero-shot role-playing abilities Njifenjou et al. (2024), prompt-driven user simulation has become the dominant approach. For example, LLMs have been used with carefully designed predefined profiles to align with human beliefs Chuang et al. (2024), simulate consultation scenarios with users exhibiting varying personalities and needs in ToD systems Zhang et al. (2024a), and model user preferences in conversational recommendation systems Yoon et al. (2024).
@@ 3 Task Definition
We formulate user simulation as a dialogue refactoring task, aiming to replicate multi-turn user behavior in target dialogues. Given a target dialogue $d_{i}=\{(u_{i1},r_{i1}),\dots,(u_{ij},r_{ij})\}$ between a user $U_{i}$ and a response model $R_{i}$, where $u_{ij}$ and $r_{ij}$ represent the $j$-th turn user utterance and the corresponding model response, respectively.
To achieve high-fidelity simulation of user responses within a given context, we aim to minimize the utterance-level distance $D_{\text{utt}}(u_{ij},u_{ij}^{\prime})$ where $u_{ij}^{\prime}\sim P_{prob}(\cdot|c_{ij}^{\prime},U_{i}^{\prime})$. Here, $c_{ij}^{\prime}$ represents the context used by the user simulator $U_{i}^{\prime}$ to generate $u_{ij}^{\prime}$. This optimization ultimately leads to reducing the dialogue-level distance, as formulated in Eq. 1.
$\min_{d_{i}^{\prime}\sim P_{prob}(\cdot|U_{i}^{\prime})}D_{\text{dia}}(d_{i},d_{i}^{\prime})$ (1)
where $D_{\text{dia}}$ represents a distance function that evaluates the user utterances between the simulated dialogue $d_{i}^{\prime}$ and the real conversation $d_{i}$.
Recent studies show that role-playing with specific user profiles ($P_{i}$) can effectively achieve diverse user simulations Liu et al. (2023). However, unlike celebrities or well-known characters, user profiles in real-world conversations are often implicit and difficult to obtain Wang et al. (2024a).
To address this, we reformulate the task by first extracting the implicit user profile from the given dialogue using profile extractor $P_{\text{extractor}}$, and then reconstructing a closer dialogue, with an emphasis on the user’s utterances, as described in Eq. 2.
$\min_{d_{i}^{\prime}\sim P_{prob}(\cdot|U_{i}^{\prime},P_{i})}D_{\text{dia}}(d_{i},d_{i}^{\prime}),$ (2)
where $P_{i}=P_{\text{extractor}}(d_{i})$.
@@ 4 Modeling User Simulator with Implicit Profiles
To accomplish this task, we propose the User Simulator with Implicit Profiles (USP) framework, as illustrated in Figure 2, which aims to minimize the objective in Eq. 2 while ensuring authenticity, consistency, and diversity.
@@ 4.1.1 User Profile Schema
The OF focuses on common topics in human conversation (Cheng et al., 2024b; Dunbar et al., 1997) including Scene-Consistent Attributes (such as age, gender, and location) and Scene-Related Attributes (such as goal,and task details). SC considers both external and internal personality dimensions represented by language style Wang et al. (2024a) and the Big-Five Traits Different from previous work Cheng et al. (2024b); Tu et al. (2024), we reformulate discrete attributes into coherent narrative descriptions to achieve greater generalization and flexibility.
@@ 4.1.2 User Profile Extractor
To obtain such a user profile, we design an LLM-driven user profile extractor extracting the implicit user profile from the human-LLM conversation. The extractor first leverages advanced LLM (such as GPT-4o) to extract the user character attributes mentioned above with a well-designed prompt. Then, the extractor collects the valid attributes (No empty) together and polishes them into natural language descriptions. Further prompt details regarding the extractor can be found in Appendix A.2.
@@ 4.1.3 Profile Quality Verification
Due to the reliance on predefined user profiles in existing role-playing approaches Zhou et al. (2024b), the correlation between user profiles and conversations has been largely overlooked. To address this limitation, we propose Dialogue Profile Consistency (DPC) for evaluating OF based on atomic fact verification that measures both precision (DP.P) and recall (DP.RJandaghi et al. (2024).
Drawing inspiration from FactScore Min et al. (2023) and Con.Score Madotto et al. (2019), we first introduce Factual Consistency (Fact.Con), as defined in Eq. 3. Given a target $T$, we evaluate the consistency between the source and target by decomposing $T$ into atomic facts $af_{k}$ using an atomic fact generator (afg). We then compute the natural language inference (NLI) score for each atomic fact with respect to the source $S$.
@@
Additionally, we use a validation score (Val.Score) to assess the quality of SC descriptions based on the dialogue, rating them on a scale from 1 to 5 using GPT-4o. Detailed prompts in Appendix D.
@@ 4.2 Conditional Supervised Fine-Tuning
To empower the LLM with the general capability to simulate diverse users at the utterance level, we utilize conditional supervised fine-tuning based on user profiles. It enables the LLM to learn the conditional generation mapping based on both the extracted profile $P_{i}$ and context $c_{ij}^{\prime}$. As subtle misalignment between the core objectives of the user simulator and the response model, the SFT language modeling loss is adjusted as shown in Eq. 4.
$l_{\text{lm}}=\sum_{j}\sum_{k}-\log P_{prob}(u_{i,j,k}^{\prime}|u_{i,j,<k}^{\prime},c_{i,j}^{\prime},P_{i})$ (4)
where $u_{ijk}$ represents the $k$-th token of the $j$-th utterance from the $i$-th user.
@@ 4.3 Reinforcement Learning with Cycle Consistency
To further enhance conversation-level consistency, we introduce Reinforcement Learning with Cycle Consistency (RLCC), which optimizes the user simulator by aligning the reflected profile, extracted from simulated dialogues, with the target profile.
In this stage, we extract the simulator profile $P_{i}^{\prime}$ from dialogues $D^{\prime}$ generated by the user simulator $U_{i}^{\prime}$ using the profile generator, based on virtual profiles sampled by our diverse profile sampler (see Section 4.4). Our goal is to maximize the semantic similarity between the target profile $P_{i}$ and the extracted reflected profile both in objective facts and subjective characteristics. The dialogue-level reward is then distributed to each user utterance within the dialogue, as shown in Eq. 5.
$r^{cc}_{i,j}=\text{sim}(P_{i},P^{\prime}_{i})$ (5)
Then, we optimize profile recall through Proximal Policy Optimization (PPO) Schulman et al. (2017) using cycle consistency as a dialogue-level reward signal. It enhances the user simulator’s self-expression in dialogues, moving beyond simple profile adherence or precision consistency.
Additionally, to prevent reward hacking, we incorporate an AI detection model as an auxiliary reward. The final reward is formulated in Eq. 6.
@@
where $r^{ai\_detect}_{i,j}=\text{AI\_detect}(u_{i,j})$ and $\lambda=0.8$ is used to emphasize the importance of cycle consistency. The AI detection model Yang et al. (2024) and profile generator are all fine-tuned based on our train dataset, details in Appendix B.1.
4.4 Diverse Profile Sampling
To generate diverse and naturalistic user profiles that reflect real-world distributions, we propose a Density Profile Sampler that preserves distribution characteristics while ensuring coverage of underrepresented cases. Profiles are first embedded into the semantic space using SimCSE Gao et al. (2021), then projected with UMAP McInnes et al. (2018) to maintain density relationships. We then estimate the underlying profile distribution using Gaussian Kernel Density Estimation (GKDE). This enables direct sampling of realistic profiles based on probability density, while generating virtual profiles during random sampling by combining OF and SC descriptions from nearest neighbors.

### v3

@@ Abstract
User simulators are crucial for replicating human interactions with dialogue systems, supporting both collaborative training and automatic evaluation, especially for large language models (LLMs). However, current role-playing methods face challenges such as a lack of utterance-level authenticity and user-level diversity, often hindered by role confusion and dependence on predefined profiles of well-known figures. In contrast, direct simulation focuses solely on text, neglecting implicit user traits like personality and conversation-level consistency. To address these issues, we introduce the User Simulator with Implicit Profiles (USP), a framework that infers implicit user profiles from human-machine interactions to simulate personalized and realistic dialogues. We first develop an LLM-driven extractor with a comprehensive profile schema, then refine the simulation using conditional supervised fine-tuning and reinforcement learning with cycle consistency, optimizing at both the utterance and conversation levels. Finally, a diverse profile sampler captures the distribution of real-world user profiles. Experimental results show that USP outperforms strong baselines in terms of authenticity and diversity while maintaining comparable consistency. Additionally, using USP to evaluate LLM on dynamic multi-turn aligns well with mainstream benchmarks, demonstrating its effectiveness in real-world applications. We open-source related resources in https://github.com/wangkevin02/USP.
@@
Kuang Wang1, Xianfei Li1, Shenghao Yang1, Li Zhou1, Feng Jiang2,1thanks: Feng Jiang is the corresponding author and Shenzhen University of Advanced Technology is the corresponding affiliation., Haizhou Li1,3 1SRIBD, School of Data Science, The Chinese University of Hong Kong, Shenzhen, Guangdong 2Artificial Intelligence Research Institute, Shenzhen University of Advanced Technology 3Department of ECE, National University of Singapore kuangwang@link.cuhk.edu.cn, jiangfeng@suat-sz.edu.cn
@@ 1 Introduction
Figure 1: Examples of different user simulators in multi-turn human-LLM interactions. "OF" and "SC" represent objective facts and subjective characteristics, respectively. highlights inconsistencies with the target profile, while indicates inauthentic user imitation.
@@
The user simulator is designed as a proxy for real users in interactions with large language models (LLMs). It can simulate a realistic user by generating the target user’s behavior or utterances based on the specified characteristics, enabling dynamic multi-turn interactions with LLMs Wan et al. (2022) and scene reproduction Wang et al. (2024b). As a result, it becomes an effective alternative Liu et al. (2023); Ferreira et al. (2024) in scenarios where real-world human-computer interaction data is difficult to obtain, especially in domains with privacy and ethical concerns, such as medical consultations Valizadeh and Parde (2022). It also helps Simulation-to-Reality (Sim2Real) applications, such as tutorial strategies, election simulations, and public opinion research Liu et al. (2024); Zhang et al. (2024b); Chuang et al. (2024).
Recent advances in LLMs have spurred the development of user simulators, improving their naturalness and utility Deng et al. (2024); Zhang et al. (2024a). Mainstream LLM-based role-playing methods Moon et al. (2024) use predefined profiles to mimic diverse user traits. However, as LLMs are typically trained to be universally polite and helpful Lu et al. (2024), they often lack utterance-level authenticity and struggle with role confusion between user simulation and their inherent assistant nature Xu et al. (2023a), as shown in Figure 1. Models like PlatoLM Kong et al. (2024) and Parrot Sun et al. (2024) address this by training on real human-LLM conversation datasets, but they focus only on text utterances, lacking dialogue control. This limits conversation-level consistency and diverse simulation without seed context. Additionally, they both fail to capture the authentic distribution of user-level diversity, crucial for analyzing group behavior.
@@
To address the issues above, we believe that a user simulator knows users’ intrinsic characters hidden in their conversations first, and then can provide a better simulation. Therefore, we treat user simulation as a dialogue reconstruction task and propose a novel framework named the User Simulator with Implicit Profile (USP). It is decomposed into implicit profile extraction to capture the user’s underlying characteristics from the target user dialogue and conditional generation based on the profile.
@@
In this framework, we first propose an LLM-driven profile extractor to extract implicit profiles from user conversations with a well-designed profile schema. Inspired by interpersonal interaction theoryKruglanski and Higgins (2013), our profile schema contains two dimensions (objective facts (OF) and subjective characteristics (SC)) with a dozen attributes to describe the user comprehensively. Different from existing works Cheng et al. (2024b); Tu et al. (2024) using attributes as profiles, we further polish the profile attributes into natural, descriptive profiles to ensure generalization.
@@
Our experiments demonstrate that USP improves semantic and stylistic similarity in reconstructed multi-turn dialogues by approximately 34% and 43% compared to the leading baseline, with reconstruction errors reduced by half, showcasing enhanced authenticity and diversity. It achieves dialogue profile consistency comparable to GPT-4o (User w/ Profile), improving multi-turn consistency by 14% while matching single-turn performance. Additionally, USP-based multi-turn dynamic evaluation of LLMs for downstream tasks aligns closely with established benchmarks, enabling finer-grained assessment of LLM performance across diverse user groups. Our key contributions are outlined below:
@@
  • We propose a framework that infers implicit user profiles, further enhanced by conditional fine-tuning and reinforcement learning with cycle consistency, to improve simulation at both the utterance and conversation levels.
  • Experiments show that USP outperforms baselines in authenticity and diversity, maintains comparable consistency, and enables effective multi-turn dynamic evaluation of LLMs.
@@ 2.1 General User Simulator
Early user simulators focused on limited action prediction using agenda-based Schatzmann et al. (2007); Schatzmann and Young (2009) and model-based methods Asri et al. (2016); Kreyssig et al. (2018), constrained by early natural language generation capabilities—for instance, generating synthetic binary preferences in conversational recommendation systems Christakopoulou et al. (2016).
@@ 2.2 Persona-based User Simulator
Since general user simulators often struggle to capture the full spectrum of diverse user needs, it leads a growing interest in persona-based personalization to improve both controllability and diversity in simulations Takanobu et al. (2020). Some researchers attempt to leverage goal generators Takanobu et al. (2020) to create diverse user goals or retrieval-based personas derived from historical data Shi et al. (2019) to guide user simulators in task-oriented dialogue (ToD) systems.
With the rise of LLMs and their strong zero-shot role-playing capabilities Njifenjou et al. (2024), prompt-driven user simulation has become the dominant paradigm. LLMs have been used to simulate users with predefined profiles Chuang et al. (2024), model diverse personalities and needs in ToD Zhang et al. (2024a), and capture user preferences in conversational recommendation Yoon et al. (2024). Our method follows this line of work, with a focus on addressing authenticity, consistency, and diversity, which remain underexplored in related studies.
@@ 3 Task Definition
We formulate user simulation as a dialogue refactoring task to replicate multi-turn user behavior in a target dialogue $d_{i}=\{(u_{i,1},a_{i,1}),\dots,(u_{i,n},a_{i,n})\}$, where $u_{i,j}$ is the $j$-th user utterance and $a_{i,j}$ is the corresponding response answer. Our goal is to achieve high utterance-level and dialogue-level fidelity. Formally, we minimize utterance-level distance $D_{\text{utt}}(u_{i,j},u_{i,j}^{\prime})$ and dialogue-level distance $D_{\text{dia}}(d_{i},d_{i}^{\prime})$, where $u_{i,j}^{\prime}$ is the simulated utterance and $d_{i}^{\prime}$ is the simulated dialogue.
Direct simulation struggles to capture personalized traits. Recent studies Deng et al. (2024); Kong et al. (2025) demonstrate that role-playing with specific user profiles ($p_{i}$) effectively enables diverse user simulations. However, unlike well-known figures, user profiles in real-world conversations are often implicit and challenging to derive Wang et al. (2024a).
To address this, we reformulate the task by extracting the implicit user profile from the dialogue using a profile extractor $P_{\text{extractor}}$, then reconstructing the target dialogue as Eq. 1.
$\min_{\begin{subarray}{c}d_{i}^{\prime}\sim P(\cdot|p_{i},\pi_{\theta})\\ u_{i,j}^{\prime}\sim P(\cdot|c_{i,j},p_{i},\pi_{\theta})\end{subarray}}\left[D_{\text{utt}}(u_{i,j},u_{i,j}^{\prime})+\alpha D_{\text{dia}}(d_{i},d_{i}^{\prime})\right],$ (1)
where $p_{i}=P_{\text{extractor}}(d_{i})$, $\pi_{\theta}$ represents the learnable parameters of the language model, and $c_{i,j}=\{(u_{i,1},a_{i,1}),\dots,(u_{i,j-1},a_{i,j-1})\}$ denotes the ground-truth context up to the $j$-th turn. The hyperparameter $\alpha$ balances the utterance-level and dialogue-level distances.
@@ 4 Modeling User Simulator with Implicit Profiles
We propose the User Simulator with Implicit Profiles (USP) framework, shown in Figure 2, to minimize the objective in Eq. 1 across four stages. First, we build a user profile extractor with a tailored schema (Section 4.1). Then, we optimize utterance-level authenticity using conditional SFT (Section 4.2), ensure conversation-level consistency via Reinforcement Learning with Cycle Consistency (RLCC) (Section 4.4), and achieve user-level diversity through corpus distribution fitting (Section 4.3).
@@ 4.1.1 User Profile Schema
The OF focuses on common topics in human conversation (Cheng et al., 2024b; Dunbar et al., 1997), including scene-consistent attributes (such as age, gender, and location) and scene-related attributes (such as the goal and task details). SC encompasses external personality dimensions, reflected in language style Wang et al. (2024a), and internal personality dimensions, captured by the Big-Five Traits Gosling et al. (2003).
Unlike prior work Cheng et al. (2024b); Tu et al. (2024) that relies on discrete attributes for user profiles, we further reformulate these attributes into coherent narrative descriptions to enhance generalization and flexibility.
@@ 4.1.2 User Profile Extractor
To obtain such a user profile, we design an LLM-driven user profile extractor that extracts the implicit user profile from the human-LLM conversation. The extractor first leverages advanced LLM (such as GPT-4o) to extract the user character attributes mentioned above with a well-designed prompt. Then, the extractor collects the valid attributes together and polishes them into natural language descriptions. Further prompt details regarding the extractor can be found in Appendix A.2.
@@ 4.1.3 Profile Quality Verification
Existing role-playing methods rely on predefined profiles and dialogues from separate sources, either extracted from novel segments or generated by LLMs Zhou et al. (2024b); Chen et al. (2023); Shao et al. (2023) without verifying alignment Wang et al. (2024a); Cheng et al. (2024b). This overlooks the correlation between profiles and dialogues, potentially hindering simulation performance Yu et al. (2024). To provide an automatic metric for evaluating this, we propose Dialogue Profile Consistency (DPC), which frames consistency as a retrieval task Jandaghi et al. (2024). DPC employs an F1-score approach, assessing consistency through atomic fact verification by measuring both precision (DP.P) and recall (DP.R).
We first introduce Factual Consistency (Fact.Con), an adaptation of FactScore Min et al. (2023) tailored for dialogue scenarios, as defined in Eq. 2. Given a target $T$, we assess its consistency with the source by decomposing $T$ into atomic facts $af_{k}$ using an atomic fact generator (afg). We then compute the natural language inference (NLI) score for each atomic fact with respect to the source $S$.
@@
where $\text{NLI}(\cdot,\cdot)$ denotes the NLI model, implemented using prompt-based GPT-4o.
@@
Additionally, we use a Validation Score (Val.Score) to evaluate SC description quality based on the dialogue, rated on a 1–5 scale using GPT-4o(prompts detailed in Appendix D).
@@ 4.2 Conditional Supervised Fine-Tuning
To empower the LLM with the general capability to simulate diverse users at the utterance level, we utilize conditional supervised fine-tuning based on user profiles. It enables the LLM to learn the conditional generation mapping based on both the extracted profile $p_{i}$ and context $c_{i,j}$. As a subtle misalignment between the core objectives of the user simulator and the response model, the SFT language modeling loss focuses on optimizing user utterance as shown in Eq. 3.
$\min_{\pi_{\theta}}\sum_{i,j,k}-\log P(u_{i,j,k}|u_{i,j,<k},c_{i,j},p_{i},\pi_{\theta})$ (3)
where $u_{i,j,k}$ denotes the $k$-th token of the $u_{i,j}$.
4.3 Diverse Profile Sampling
We propose Diverse Profile Sampling to generate naturalistic user profiles that reflect real-world characteristic distributions. Our method first embeds constructed profiles into a semantic space using SimCSE Gao et al. (2021), followed by dimensionality reduction via UMAP McInnes et al. (2018). We then apply Gaussian Kernel Density Estimation (GKDE) to fit the underlying distribution, allowing probabilistic sampling of realistic profiles for downstream tasks such as majority representation. To further enhance diversity, we synthesize virtual profiles by combining OF and SC descriptions from nearest neighbors, producing novel yet plausible profile variants.
@@ 4.4 Reinforcement Learning with Cycle Consistency
The conditional SFT stage enables the user simulator to generate human-like utterances focused on forward consistency, producing precise responses aligned with the target profile. However, it does not ensure full reflection of the profile, i.e., backward consistency or profile recall. To overcome this and improve conversation-level consistency, we introduce Reinforcement Learning with Cycle Consistency (RLCC), which enhances alignment between the user simulator’s actual behavior reflected in simulated dialogues and the target behavior defined by the profile, ensuring a closer match to the intended target profile.
In this stage, the user simulator $u_{i}^{\prime}$ interacts with a response LLM based on a target profile $p_{i}$, sampled via Diverse Profile Sampling, to generate a simulated dialogue $d_{i}^{\prime}$. The dialogue ends when it reaches the maximum context length or a predefined turn limit (set to 10). The reflected profile $p_{i}^{\prime}$ is then extracted from $d_{i}^{\prime}$ using the profile generator. Our goal is to maximize the semantic similarity between the target profile $p_{i}$ and the reflected profile $p_{i}^{\prime}$, both in objective facts and subjective characteristics, as defined in Eq. 4.
$\max_{\pi_{\theta}}\;\mathbb{E}_{p_{i}\sim D,d_{i}^{\prime}\sim\pi_{\theta}(p_{i})}\left[\text{Sim}(p_{i},P_{extractor}(d_{i}^{\prime}))\right]$ (4)
where $\text{Sim}(\cdot,\cdot)$ is a similarity model (SimCSE Gao et al. (2021)), and $D$ denotes the virtual profiles dataset from the sampler in Section 4.3. The dialogue-level reward is uniformly attributed to each user utterance, defined as $r_{i,j}^{cc}=\text{Sim}(p_{i},P_{extractor}(d_{i}^{\prime}))$.
To prevent reward hacking, the AI detection reward is included as an auxiliary component. The final reward, defined in Eq. 5, is utilized to optimize profile recall via Proximal Policy Optimization (PPO)Schulman et al. (2017).
@@
where $r^{ai\_detect}_{i,j}=\text{AI\_detect}(u_{i,j}^{\prime})$, and $\lambda=0.8$ prioritizes cycle consistency. The AI_detect refers to a binary AI detection model Yang et al. (2024) that predicts the probability of an utterance being AI-generated. Both the AI detection model and profile generator are fine-tuned on our training dataset, with details provided in Appendix B.1.